OpenStudy (anonymous):

The height h in feet of a ball t seconds after being tossed upwards is given by the function h(t) = 84 t - 16 t^2. a. after how many seconds will it hit the ground? b. what is its maximum height?

OpenStudy (campbell_st):

well let the height = 0 and solve so you are looking at \[0 = 84t - 16t^2\] the equation can be solved by factoring

OpenStudy (campbell_st):

the max height occurs when the time t is half way between the zeros for height, found in (a) when you get the value of t, substitute it into the original equation to find the max height

OpenStudy (anonymous):

can you show me how? @campbell_st

OpenStudy (campbell_st):

well can you factor the quadratic...?

OpenStudy (campbell_st):

what if I suggested you take 4t out as a common factor, what would you have..?

OpenStudy (anonymous):

4t(21-4t) is that right? @campbell_st

OpenStudy (campbell_st):

yep thats its... so now what values of t make the equation equal zero. so you need to solve 4t = 0 and 21 - 4t = 0 what do you get

OpenStudy (anonymous):

t=0 t=-21/4 is that right?

OpenStudy (campbell_st):

thats almost correct... so the ball starts on the ground, t = 0 seconds, and then returns to the ground after t = 21/4 seconds does that make sense... you had a slight error 21 - 4t = 0..... so 21 = 4t ...etc

OpenStudy (campbell_st):

the curve is a parabola  find the middle value, substitute it to find the max height. unless you have to use calculus